# Towards Federated Graph Learning in One-shot Communication

### v6

@@ 5.2.1 Performance Comparison
The performance under the Louvain partition with 10 clients is shown in Table 1, and results with more clients and results under the Metis partition are shown in the Appendix A.3.1. Our method consistently outperforms on both accuracy and F1-macro, demonstrating the superiority of our method in deriving models with better personalization and local generalization. FedTAD encounters the OOM issue due to its graph diffusion operation, and FedLoG encounters the OOM issue due to its -hop subgraph construction for each node. FedSpray and FedLoG also focus on minor and underrepresented classes, but they require numerous communications for optimization under non-IID scenarios, offering little advantage in one-shot communication. Furthermore, our method (1 communication round) even outperforms other multi-round federated learning methods (100 communication rounds) on average, which is shown in the Appendix A.3.2. More results on Computers, Photo, CS, and Physics datasets are listed in the Appendix A.3.6. We also include the experiments with other prevalent graph models (GraphSage [19], GAT [45], and SGC [49]) in Appendix A.3.7.
5.2.2 Ablation Study
@@
We conduct the ablation study to analyze our method. We compare our method with 4 variants: (1) ft: only fine-tuning to obtain on local graph data, (2) ft(focal): fine-tuning with focal loss [39] to handle data imbalance, (3) ft(fixed ): in the second stage, performing distillation with a fixed for all nodes; (4) ours w.o. HRE: performing node adaptive distillation along with fine-tuning, without HRE, and (5) ours: performing the complete method (including HRE).
@@
The results under the Louvain partition and Metis partition are shown in Figure 4 and Figure 4, respectively. The focal loss cannot effectively address the non-IID graph data since it’s hard to determine the proper weights. Fixed cannot balance improvements in accuracy and F1-macro as it cannot distinguish the majority from the minority. For nodes in major classes, the global knowledge of may have a negative impact, as demonstrated by the left sub-figure in Figure 5. This motivates us to develop node-adaptive distillation, which is more flexible to determine when to introduce global knowledge of in fine-tuning. HRE strategy further improves performance by generating a better global pseudo-graph with more precise estimations. Furthermore, we justify the rationale for considering homophily in determining the major classes by values in Eq. 2 in the right panel of Figure 5. The detailed analysis is in the Appendix A.3.3.
@@
| Method | Cora | CiteSeer | ogbn-arxiv | |||
| Acc. | F1 | Acc. | F1 | Acc. | F1 | |
| O-pFGL (Ours) | 75.04 0.50 | 57.26 1.93 | 70.87 0.21 | 56.67 0.20 | 70.98 0.09 | 36.41 0.43 |
5.2.3 More Experiments
Inductive Performance.
Model Heterogeneity.
We conduct experiments on 10 clients with heterogeneous models and configurations are described in Appendix A.2. Under model heterogeneity, most pFGL methods do not work since they rely on the model parameter aggregation. The results under Louvain and Metis partitions are presented in Table 3 and Table 11, respectively. The results demonstrate the superior performance of our method in both accuracy and F1-macro. Details are in the Appendix A.3.4.
Hyper-parameters Analysis.
We study our method’s sensitivity to the hyper-parameters in Appendix A.3.8. Our method is generally robust under a proper range of hyper-parameters values.
5.2.4 Communication Analysis
We analyze the communication costs both theoretically and empirically. Our method involves uploading the class-wise statistics () and downloading the generated global pseudo-graph (, is the pre-set number of nodes in the global pseudo-graph, which is small as described in Appendix A.2). Both are independent of the model size. Figure 6 reports the communication costs of baseline methods and our method on the Cora datasets under the Louvain and Metis partition. The results show that our method has lower communication costs. We include the detailed theoretical analysis of communication, computation, and memory complexity in the Appendix A.4. We also introduce an equivalent server-efficient variant of our method in the Appendix A.5.
6 Conclusion
We propose the first one-shot personalized federated graph learning method that supports model heterogeneity and Secure Aggregation. Specifically, our method first estimates class-wise feature distribution statistics on each client, augmented by our proposed HRE, and then aggregates these statistics on the server to generate the global pseudo-graph. To further enhance personalization and generalization, we propose a novel two-stage personalized training with node-adaptive distillation, which balances local information and global insights. Extensive experiments demonstrate the superior performance of our method across various settings.
Appendix A Appendix
A.1 Illustration for Determining in Node Adaptive Distillation
We illustrate how to determine for each node in each client’s local graph. We pre-compute the vector by Eq. 3 and soft labels by Label Propagation in Section 4.2. Considering soft label of each node, if a node is predicted to class with a small (e.g. node in Figure 7), it likely belongs to major classes with high homophily. In this case, is small, and primarily learns through supervised fine-tuning; If a node is predicted to class with a large or has a blended prediction (e.g. node in Figure 7), it likely belongs to minor classes or has low homophily. These nodes are often neglected during supervised fine-tuning. Thus should be large to incorporate the corresponding global knowledge from .
A.2 Experimental Setup
| Dataset | Task | #Nodes | #Edges | #Features | #Classes | Train/Val/Test | Description |
| Cora | Transductive | 2,708 | 10,556 | 1433 | 7 | 20%/40%/40% | Citation Network |
| CiteSeer | Transductive | 3,327 | 4,732 | 3703 | 6 | 20%/40%/40% | Citation Network |
| PubMed | Transductive | 19,717 | 44,338 | 500 | 3 | 20%/40%/40% | Citation Network |
| Ogbn-arxiv | Transductive | 169,343 | 1,166,243 | 128 | 40 | 60%/20%/20% | Citation Network |
| Ogbn-products | Transductive | 2,449,029 | 61,859,140 | 100 | 47 | 10%/5%/85% | Co-purchase Network |
| Computers | Transductive | 13,381 | 245,778 | 767 | 10 | 20%/40%/40% | Co-purchase Network |
| Photo | Transductive | 7,487 | 119,043 | 745 | 8 | 20%/40%/40% | Co-purchase Network |
| CS | Transductive | 18,333 | 81,894 | 6,805 | 15 | 20%/40%/40% | Co-author Network |
| Physics | Transductive | 34,493 | 247,962 | 8,415 | 5 | 20%/40%/40% | Co-author Network |
| Flickr | Inductive | 89,250 | 899,756 | 500 | 7 | 60%/20%/20% | Image Network |
| Inductive | 232,965 | 114,615,892 | 602 | 41 | 80%/10%/10% | Social Network | |
| Reddit2 | Inductive | 232,965 | 23,213,838 | 602 | 41 | 65%/10%/25% | Social Network |
Dataset Description.
We conduct experiments on 12 graph datasets, and the summary is listed in Table 4. Our selected datasets cover a wide range from 1) small-scale to large-scale, 2) transductive to inductive, and 3) academic domain to industry domain.
Hyper-parameters and Model Configurations.
In our one-shot personalized federated graph learning setting, we limit the communication round of the 5 pFGL methods and the FedAvg method to a single round. For 2 OFL methods (DENSE and Co-Boost), the generators are trained to generate node features, and the topology structure is constructed using the -Nearest Neighbors strategy as outlined in [69]. Note that we perform additional fine-tuning processes on FedAvg, FedTAD, DENSE, and Co-Boost to ensure better performance.
For each client, we employ a two-layer GCN [31] with the hidden layer dimension set to 64 by default. For baseline methods, the number of local training epochs is tuned from 1 to 100, and hyper-parameters are set as recommended in their respective papers or fine-tuned for optimal overall performance. For our method, we keep the size of the generated global pseudo-graph as small as possible. On small-scale Cora, CiteSeer, and PubMed datasets, we set the number of nodes in each class in the global pseudo-graph to 1. For other datasets, we set the number of nodes in each class to percent of the total number of nodes of the corresponding class in the training set. We set as 0.25%, 0.04%, 1%, 1%, 1%, 0.5%, 0.2%, 0.05%, 0.05% on ogbn-arxiv, ogbn-products, Computers, Photo, CS, Physics, Flickr, Reddit, and Reddit2 datasets. We use the Adam optimizer and tune the learning rate. We tune within [0.95, 1] and tune according to the scale of the graphs. We tune the within [0.1, 1] to control the range of . For other hyper-parameters in global pseudo-graph generation, we adopt the setting in [51].
Under model heterogeneity, we adopt 10 different GNN models categorized into two types: coupled GNNs and decoupled GNNs. For coupled GNNs, we adopt 6 models, which are: 2-layer GCN with 64, 128, 256, and 512 hidden dimensions, and 3-layer GCN with 64 and 512 hidden dimensions. For decoupled GNNs, we adopt APPNP [16], SSGC [68], SGC (=2) and SGC (=4) [49].
Experimental Environments.
The experimental machine is with Intel(R) Xeon(R) CPU E5-2680 v4@2.40GHz, and 6NVIDIA GeForce RTX 3090. The operating system is Ubuntu 20.04, and the version of PyTorch is 2.1.2. We use PyG 2.5.2 [11].
| Method | Cora | CiteSeer | PubMed | ogbn-arxiv | ogbn-products | |||||
| Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | |
| O-pFGL (Ours) | 69.84 1.76 | 49.37 1.98 | 68.69 0.04 | 50.26 0.80 | 82.04 0.20 | 58.79 0.68 | 70.61 0.05 | 32.28 0.68 | 87.03 0.00 | 21.87 0.05 |
A.3 More Experimental Results and Detailed Analysis
A.3.1 Performance Comparison
We first conduct experiments in the transductive learning setting and report the accuracy and F1-macro metrics. The datasets range from small-scale graphs to large-scale graphs with millions of nodes. We also vary the number of clients to evaluate the stability of methods. With 10 clients, the experimental results under Louvain and Metis partitions are shown in Table 1 and Table 6, respectively. With 20 clients, the experimental results under Louvain and Metis partitions are shown in Table 5 and Table 8, respectively. Our method consistently outperforms others in terms of accuracy and F1-macro. Under the non-IID scenarios, pFGL methods show marginal or even negative performance gain over Standalone, highlighting the ineffectiveness of current pFGL methods in personalizing models on non-IID graph data in a single communication round. The state-of-the-art OFL methods (DENSE and Co-Boost) also have difficulty generating high-quality pseudo-data by the ensemble model when each model is trained on non-IID local graph data. FedTAD encounters OOM when handling million-scale graphs due to its graph diffusion operation. FedLoG encounters OOM issues due to its node-wise -hop subgraph construction. In contrast, our method remains scalable. FedGTA performs badly because its aggregation is based on the unreliable local models’ predictions under non-IID data in a single round. FedSpray focuses on minor classes, but it requires numerous communications for optimization, offering little advantage over others in one-shot communication. FedLoG focuses on local generalization, but it cannot work well under severe non-IID scenarios within one-shot communication.
| Method | Cora | CiteSeer | PubMed | ogbn-arxiv | ogbn-products | |||||
| Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | |
| O-pFGL (Ours) | 81.79 0.23 | 50.85 1.27 | 72.76 0.24 | 50.94 0.80 | 83.99 0.18 | 61.03 1.21 | 71.06 0.09 | 41.35 0.37 | 86.90 0.01 | 28.90 0.05 |
We also compare our method with ReNode [6], a representative graph imbalance learning method, which aims to train robust graph models under imbalanced graph data. The comparison results are shown in Table 7. The results show that ReNode cannot effectively model the imbalanced graph data in our scenarios.
| Louvain | Cora | CiteSeer | PubMed | |||
| Acc. | F1 | Acc. | F1 | Acc. | F1 | |
| O-pFGL (Ours) | 76.43 1.24 | 61.58 2.16 | 71.61 0.40 | 58.24 0.96 | 82.71 0.03 | 66.40 0.34 |
| Metis | Cora | CiteSeer | PubMed | |||
| Acc. | F1 | Acc. | F1 | Acc. | F1 | |
| O-pFGL (Ours) | 81.79 0.23 | 50.85 1.27 | 72.76 0.24 | 50.94 0.80 | 83.99 0.18 | 61.03 1.21 |
| Method | Cora | CiteSeer | PubMed | ogbn-arxiv | ogbn-products | |||||
| Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | |
| O-pFGL (Ours) | 78.31 0.88 | 37.21 1.96 | 68.88 0.22 | 41.95 0.25 | 83.94 0.05 | 56.61 1.36 | 70.87 0.18 | 36.06 0.43 | 87.15 0.02 | 23.41 0.26 |
A.3.2 Compared with Multi-Rounds FL Methods
To demonstrate the superiority of our method, we further compare our method with multi-round federated learning methods, including FedAvg, FedPUB, FedGTA, FedTAD, FedSpray, and FedLoG. We set the communication round of these 6 methods to 100. The experimental results on 5 multi-scale graph datasets under Louvain and Metis partitions are shown in Table 9 and Table 10, respectively. Our method still outperforms in most cases and remains the best average performance.
| Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | |
| O-pFGL (Ours) | 76.43 1.24 | 61.58 2.16 | 71.61 0.40 | 58.24 0.96 | 82.71 0.03 | 66.40 0.34 | 70.93 0.59 | 36.66 0.22 | 86.63 0.02 | 27.21 0.15 | 77.66 | 50.02 |
| Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | |
| O-pFGL (Ours) | 81.79 0.23 | 50.85 1.27 | 72.76 0.24 | 50.94 0.80 | 83.99 0.18 | 61.03 1.21 | 71.06 0.09 | 41.35 0.37 | 86.90 0.01 | 28.90 0.05 | 79.30 | 46.61 |
A.3.3 Detailed Ablation Study
We conduct the ablation study to evaluate the performance gain of our method. We compare our method with 4 variants: (1) ft: only fine-tuning to obtain on local graph data, (2) ft(focal): fine-tuning with focal loss [39] to handle data imbalance, (3) ft(fixed ): in the second stage, performing distillation with a fixed for all nodes; (4) ours w.o. HRE: performing node adaptive distillation along with fine-tuning, without HRE, and (5) ours: performing the complete method (including HRE).
Figure 4 and Figure 4 show our experimental results of accuracy and F1-macro under Louvain and Metis partition, respectively. We make the following 4 observations. (O1) Comparing ft and ft(focal), the focal loss in ft(focal) cannot effectively address the non-IID graph data. We believe the reasons are that it’s hard to determine the proper weights under extreme non-IID scenarios. (O2) Comparing ft and ft(fixed ), a fixed in distillation cannot balance the improvements of accuracy and F1-macro. We believe the reason is that it does not distinguish the majority and minority in the distillation. For nodes in major classes with high homophily, the global knowledge of may have a negative impact, as demonstrated by the left panel in Figure 5. (O3) Comparing ft and ours w.o. HRE, the node adaptive distillation is more flexible in determining when to introduce the global knowledge of in fine-tuning. Thus ours w.o. HRE achieves better accuracy and F1-macro. (O4) Comparing ours w.o. HRE and ours, the HRE strategy further improves performance by generating a higher-quality global pseudo-graph through more precise estimations.
A.3.4 Performance under Model Heterogeneity
We conduct experiments on 10 clients with heterogeneous models. The model configuration is detailed in Section A.2. Under model heterogeneity, most pFGL methods (except FedTAD and FedSpray) do not work since they rely on the model parameter aggregation. Thus, we compare our method with FedTAD, FedSpray, DENSE, and Co-Boost. The results under Louvain and Metis partitions are presented in Table 3 and Table 11. The results demonstrate the superior performance of our method in both accuracy and F1-macro.
| Method | Cora | CiteSeer | ogbn-arxiv | |||
| Acc. | F1 | Acc. | F1 | Acc. | F1 | |
| O-pFGL (Ours) | 79.64 0.94 | 44.86 2.04 | 71.82 0.34 | 47.57 0.98 | 70.31 0.06 | 40.31 0.11 |

### v7

@@ Introduction
We conduct experiments on various datasets with different properties to compare the performance. (1) Our method consistently outperforms on both accuracy and F1-macro, demonstrating the superiority of our method in deriving models with better personalization and generalization. As shown in Table 1, most baselines are ineffective in one-shot communication under data heterogeneity. FedSpray and FedLoG focus on underrepresented classes, but they require numerous communications for optimization, offering little advantage in one-shot communication. FedGM and GHOST fail to handle severe heterogeneity, as biased gradients and skewed local data lead to inferior model performance. And they do not address the personalization issues. FedTAD, FedLOG, FedGM, and GHOST are not scalable and face out-of-memory problems on large-scale datasets. (2) The superiority of our method extends to heterophilic graph datasets, which is shown in Table 2. (3) Furthermore, our one-shot method achieves better average performance than multi-round federated methods that run for 100 rounds, which is shown in Appendix D.2. Full results for more clients, Metis partition, more datasets, and other prevalent GNN models (GraphSage, GAT, and SGC) are available in Appendix D.1, Appendix D.7, and Appendix D.8.
Model Heterogeneity
We conduct experiments with heterogeneous client models. The results are available in Appendix D.5. Our method outperforms other methods.
Inductive Performance
We conduct experiments on inductive datasets, and the results are available in Appendix D.6. Our method is scalable and outperforms.
Ablation Study
@@
We compare our method with 4 variants: (1) ft: only fine-tuning to obtain on local graph data, (2) ft(focal): fine-tuning with focal loss (Ross and Dollár 2017) to handle data imbalance, (3) ft(fixed ): in the second stage, performing distillation with a fixed for all nodes; (4) ours w.o. HRE: performing our method without HRE.
@@
The results are shown in Figure 3. (1) The focal loss is ineffective on non-IID graph data since it’s hard to determine proper weights. (2) Fixed cannot balance improvements in accuracy and F1-macro as it cannot distinguish majority from minority. For major classes, the global knowledge of may have a negative impact, as shown in the left of Figure 5. This motivates our node adaptive distillation, which is more flexible in determining when to introduce global knowledge in fine-tuning. (3) HRE further improves performance via more precise estimations on homophilic graphs. Furthermore, we justify the rationale of considering homophily in determining the major classes by in Eq. 2 in the right of Figure 5. More analysis is in Appendix D.3.
Hyper-parameters Analysis
We study our method’s sensitivity to the hyper-parameters in Appendix E. Our method is robust under a proper range of hyper-parameters values. We also provide guidance on selecting proper values.
Communication Costs
We analyze communication costs both theoretically and empirically. Our method involves uploading class-wise statistics () and downloading the generated global surrogate graph (). Here is a small number detailed in Appendix C. Both are independent of the model size. As empirically demonstrated on the Cora dataset in Figure 4, our method achieves the lowest communication cost. Furthermore, we introduce a server-efficient variant in Appendix G to minimize server-side overhead.
Complexity Analysis
Additional complexity of our method comes from feature distribution estimation and surrogate graph generation. The former is extremely lightweight, requiring only a single forward pass, and can be offloaded to CPUs to ensure scalability. The surrogate graph generation is also scalable as the optimization is confined to a small, fixed-size graph (e.g., under 300 nodes in our experiments). This process is equivalent to forward and backward on this small graph. In our experiments, our optimization for the surrogate graph takes less than 2 minutes on the RTX 3090 GPU across all datasets. For comparison on the ogbn-arxiv dataset, FedTAD requires 6 minutes to calculate its diffusion matrix, GHOST needs 12 minutes to compute its Topology Consistency Criterion, and FedGM takes over 20 minutes for local graph synthesis. A formal theoretical complexity analysis is provided in Appendix F.
Conclusion
In this paper, we introduce O-pFGL, the first one-shot personalized federated graph learning method for heterogeneous clients. It effectively balances personalization and generalization while supporting model heterogeneity and Secure Aggregation. In O-pFGL, clients first estimate class-wise feature statistics, which the server aggregates to generate a global surrogate graph without compromising raw data. Second, we introduce a novel two-stage personalized training with node-adaptive distillation, which effectively balances local knowledge with global insights. Extensive experiments demonstrate the state-of-the-art performance of our method across diverse and challenging settings.
Appendix A Illustration for Determining in Node Adaptive Distillation
We illustrate how to determine for each node in each client’s local graph. We pre-compute the vector by Eq. 16 and soft labels by Label Propagation in Section Feature Distribution Estimation. Considering the soft label of each node, if a node is predicted to class with a small (e.g. node in Figure 6), it likely belongs to major classes with high homophily. In this case, is small, and primarily learns through supervised fine-tuning; If a node is predicted to class with a large or has a blended prediction (e.g. node in Figure 6), it likely belongs to minor classes or has low homophily. These nodes are often neglected during supervised fine-tuning. Thus should be large to incorporate the corresponding global knowledge from .
Appendix B Compatibility with Classical Secure Aggregation Implementation
The uploading and aggregation process of our method is compatible with the classical Secure Aggregation implementation since it can be decomposed into a series of weighted average operations. We show the pseudo-code of a classical implementation in Algorithm 1.
Input: A set of clients . Each client has local data to compute .
Output: .
Appendix C Experimental Setup
| Dataset | Task | Nodes | Edges | Features | Classes | Train/Val/Test | Description | Property |
| Cora | Transductive | 2,708 | 10,556 | 1433 | 7 | 20%/40%/40% | Citation Network | Homophilic |
| CiteSeer | Transductive | 3,327 | 4,732 | 3703 | 6 | 20%/40%/40% | Citation Network | Homophilic |
| PubMed | Transductive | 19,717 | 44,338 | 500 | 3 | 20%/40%/40% | Citation Network | Homophilic |
| Ogbn-arxiv | Transductive | 169,343 | 1,166,243 | 128 | 40 | 60%/20%/20% | Citation Network | Homophilic |
| Ogbn-products | Transductive | 2,449,029 | 61,859,140 | 100 | 47 | 10%/5%/85% | Co-purchase Network | Homophilic |
| Computers | Transductive | 13,381 | 245,778 | 767 | 10 | 20%/40%/40% | Co-purchase Network | Homophilic |
| Photo | Transductive | 7,487 | 119,043 | 745 | 8 | 20%/40%/40% | Co-purchase Network | Homophilic |
| CS | Transductive | 18,333 | 81,894 | 6,805 | 15 | 20%/40%/40% | Co-author Network | Homophilic |
| Physics | Transductive | 34,493 | 247,962 | 8,415 | 5 | 20%/40%/40% | Co-author Network | Homophilic |
| Actor | Transductive | 7,600 | 30,019 | 932 | 5 | 50%/25%/25% | Actor Network | Heterophlic |
| genius | Transductive | 421,961 | 984,979 | 12 | 2 | 50%/25%/25% | Actor Network | Heterophlic |
| Flickr | Inductive | 89,250 | 899,756 | 500 | 7 | 60%/20%/20% | Image Network | Heterophilic |
| Inductive | 232,965 | 114,615,892 | 602 | 41 | 80%/10%/10% | Social Network | Homophilic | |
| Reddit2 | Inductive | 232,965 | 23,213,838 | 602 | 41 | 65%/10%/25% | Social Network | Homophilic |
Dataset Description.
We conduct experiments on 14 diverse real-world graph datasets, and the summary is listed in Table 3. Our datasets cover a wide range from 1) small-scale to large-scale, 2) transductive to inductive, 3) homophilic to heterophilic, and 4) academic domain to industry domain.
Hyper-parameters and Model Configurations.
In our one-shot personalized federated graph learning setting, we limit the communication round of the 7 pFGL methods and the FedAvg method to a single round. For 2 OFL methods (DENSE and Co-Boost), the generators are trained to generate node features, and the topology structure is constructed using the -Nearest Neighbors strategy as outlined in (Zhu et al. 2024b). Note that we perform additional fine-tuning processes on FedAvg, FedTAD, FedGM, GHOST, DENSE, and Co-Boost to ensure better performance.
For each client, we employ a two-layer GCN (Kipf and Welling 2016) with the hidden layer dimension set to 64 by default. For baseline methods, the number of local training epochs is tuned from 1 to 100, and hyper-parameters are set as recommended in their respective papers or fine-tuned for optimal overall performance. For our method, we keep the size of the generated global surrogate graph as small as possible. On small-scale Cora, CiteSeer, and PubMed datasets, we set the number of nodes in each class in the global surrogate graph to 1. For other datasets, we set the number of nodes in each class to percent of the total number of nodes of the corresponding class in the training set. We set as 0.25%, 0.04%, 1%, 1%, 1%, 0.5%, 0.2%, 0.05%, 0.05%, 2%, 0.1% on ogbn-arxiv, ogbn-products, Computers, Photo, CS, Physics, Flickr, Reddit, Reddit2, Actor, and genius datasets. We use the Adam optimizer and tune the learning rate. We tune within [0.95, 1] and tune according to the scale of the graphs. We tune the within [0.1, 1] to control the range of . For other hyper-parameters in global surrogate graph generation, we adopt the setting in (Xiao et al. 2024).
Under model heterogeneity, we adopt 10 different GNN models categorized into two types: coupled GNNs and decoupled GNNs. For coupled GNNs, we adopt 6 models, which are: 2-layer GCN with 64, 128, 256, and 512 hidden dimensions, and 3-layer GCN with 64 and 512 hidden dimensions. For decoupled GNNs, we adopt APPNP (Gasteiger, Bojchevski, and Günnemann 2018), SSGC (Zhu and Koniusz 2021), SGC (=2) and SGC (=4) (Wu et al. 2019).
Experimental Environments.
The experimental machine is equipped Intel(R) Xeon(R) CPU E5-2680 v4@2.40GHz, and 4NVIDIA GeForce RTX 3090. The operating system is Ubuntu 20.04, and the version of PyTorch is 2.1.2. We use PyG 2.5.2 in the implementation.
| Method | Cora | CiteSeer | PubMed | ogbn-arxiv | ogbn-products | |||||
| Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | |
| FedGM | 60.680.43 | 28.520.26 | 58.240.29 | 35.080.27 | 72.610.12 | 31.290.22 | 48.920.39 | 1.820.05 | OOM | |
| GHOST | 62.660.20 | 31.220.35 | 62.780.23 | 43.720.37 | 81.790.05 | 58.780.23 | 67.810.29 | 16.750.79 | OOM | |
| O-pFGL | 69.841.76 | 49.371.98 | 68.690.04 | 50.260.80 | 82.040.20 | 58.790.68 | 70.610.05 | 32.280.68 | 87.030.00 | 21.870.05 |
Appendix D More Experimental Results and Detailed Analysis
D.1 Performance Comparison
We first conduct experiments in the transductive learning setting and report the accuracy and F1-macro metrics. The datasets range from small-scale graphs to large-scale graphs with millions of nodes. We also vary the number of clients to evaluate the stability of methods. With 10 clients, the experimental results under Louvain and Metis partitions are shown in Table 1 and Table 5, respectively. With 20 clients, the experimental results under Louvain and Metis partitions are shown in Table 4 and Table 7, respectively. Our method consistently outperforms others in terms of accuracy and F1-macro. Under the non-IID scenarios, pFGL methods show marginal or even negative performance gain over Standalone, highlighting the ineffectiveness of current pFGL methods in personalizing models on non-IID graph data in a single communication round. The state-of-the-art OFL methods (DENSE and Co-Boost) also have difficulty generating high-quality pseudo-data by the ensemble model when each model is trained on non-IID local graph data. FedTAD encounters OOM when handling million-scale graphs due to its graph diffusion operation. FedLoG encounters OOM issues due to its node-wise -hop subgraph construction. In contrast, our method remains scalable. FedGTA performs badly because its aggregation is based on the unreliable local models’ predictions under non-IID data in a single round. FedSpray focuses on minor classes, but it requires numerous communications for optimization, offering little advantage over others in one-shot communication. FedLoG focuses on local generalization, but it cannot work well under severe non-IID scenarios within one-shot communication. FedGM synthesized graphs with gradient matching locally. But with biased local graph data, the gradients could also be skewed, which limits the effectiveness under heterogeneous client graph data. GHOST integrates aligned proxy models. However, the proxy models are aligned with skewed client graph data. Also, it does not address the personalized issues under non-IID graph data, leading to inferior performance under severe data non-IID scenarios. Moreover, FedGM and GHOST are not scalable. On large-scale datasets, the gradient matching consumes more GPU memory. GHOST needs to reconstruct a pseudo-graph with the same size as the local graph by feature matrix multiplication, which is of high complexity.
| Method | Cora | CiteSeer | PubMed | ogbn-arxiv | ogbn-products | |||||
| Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | |
| FedGM | 75.410.11 | 23.431.18 | 65.810.37 | 36.670.42 | 76.410.37 | 29.840.43 | 45.481.04 | 1.910.27 | OOM | |
| GHOST | 76.630.07 | 31.680.05 | 64.920.07 | 37.490.12 | 83.560.04 | 52.320.96 | 68.220.22 | 22.430.48 | OOM | |
| O-pFGL | 81.790.23 | 50.851.27 | 72.760.24 | 50.940.80 | 83.990.18 | 61.031.21 | 71.060.09 | 41.350.37 | 86.900.01 | 28.900.05 |
We also compare our method with ReNode (Chen et al. 2021a), a representative graph imbalance learning method, which aims to train robust graph models under imbalanced graph data. The comparison results are shown in Table 6. The results show that ReNode cannot effectively model the imbalanced graph data in our scenarios.
| Louvain | Cora | CiteSeer | PubMed | |||
| Acc. | F1 | Acc. | F1 | Acc. | F1 | |
| O-pFGL | 76.431.24 | 61.582.16 | 71.610.40 | 58.240.96 | 82.710.03 | 66.400.34 |
| Metis | Cora | CiteSeer | PubMed | |||
| Acc. | F1 | Acc. | F1 | Acc. | F1 | |
| O-pFGL | 81.790.23 | 50.851.27 | 72.760.24 | 50.940.80 | 83.990.18 | 61.031.21 |
| Method | Cora | CiteSeer | PubMed | ogbn-arxiv | ogbn-products | |||||
| Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | |
| FedGM | 73.200.24 | 18.970.20 | 62.100.14 | 28.690.02 | 77.500.32 | 31.060.30 | 49.130.35 | 1.690.06 | OOM | |
| GHOST | 73.550.37 | 18.780.31 | 62.600.40 | 28.440.20 | 77.420.47 | 31.170.52 | 69.290.07 | 29.840.34 | OOM | |
| O-pFGL | 78.310.88 | 37.211.96 | 68.880.22 | 41.950.25 | 83.940.05 | 56.611.36 | 70.870.18 | 36.060.43 | 87.150.02 | 23.410.26 |
D.2 Compared with Multi-Rounds FL Methods
To demonstrate the superiority of our method, we further compare our method with multi-round federated learning methods, including FedAvg, FedPUB, FedGTA, FedTAD, FedSpray, FedLoG, and FedGM. We set the communication round of these 6 methods to 100. The experimental results on 5 multi-scale graph datasets under Louvain and Metis partitions are shown in Table 8 and Table 9, respectively. Our method still outperforms in most cases and remains the best average performance.
| Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | |
| FedGM | 69.390.56 | 44.040.55 | 61.700.24 | 42.380.45 | 79.300.29 | 44.690.39 | 54.650.69 | 8.650.60 | OOM | N/A | N/A | |
| O-pFGL | 76.431.24 | 61.582.16 | 71.610.40 | 58.240.96 | 82.710.03 | 66.400.34 | 70.930.59 | 36.660.22 | 86.630.02 | 27.210.15 | 77.66 | 50.02 |
| Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | |
| FedGM | 76.920.22 | 31.660.40 | 65.810.37 | 36.670.42 | 79.860.31 | 36.730.70 | 50.150.48 | 8.550.74 | OOM | N/A | N/A | |
| O-pFGL | 81.790.23 | 50.851.27 | 72.760.24 | 50.940.80 | 83.990.18 | 61.031.21 | 71.060.09 | 41.350.37 | 86.900.01 | 28.900.05 | 79.30 | 46.61 |
D.3 Detailed Ablation Study
We conduct the ablation study to evaluate the performance gain of our method. We compare our method with 4 variants: (1) ft: only fine-tuning to obtain on local graph data, (2) ft(focal): fine-tuning with focal loss (Ross and Dollár 2017) to handle data imbalance, (3) ft(fixed ): in the second stage, performing distillation with a fixed for all nodes; (4) ours w.o. HRE: performing node adaptive distillation along with fine-tuning, without HRE.
Figure 3 and Figure 7 show our experimental results of accuracy and F1-macro under Louvain and Metis partition, respectively. We make the following 4 observations. (O1) Comparing ft and ft(focal), the focal loss in ft(focal) cannot effectively address the non-IID graph data. We believe the reasons are that it’s hard to determine the proper weights under extreme non-IID scenarios. (O2) Comparing ft and ft(fixed ), a fixed in distillation cannot balance the improvements of accuracy and F1-macro. We believe the reason is that it does not distinguish the majority and minority in the distillation. For nodes in major classes with high homophily, the global knowledge of may have a negative impact, as demonstrated by the left panel in Figure 5. (O3) Comparing ft and ours w.o. HRE, the node adaptive distillation is more flexible in determining when to introduce the global knowledge of in fine-tuning. Thus ours w.o. HRE achieves better accuracy and F1-macro. (O4) Comparing ours w.o. HRE and ours, the HRE strategy further improves performance by generating a higher-quality global surrogate graph through more precise estimations.
D.4 Different Adjacency Matrix Implementations
In our method, we derive the adjacency matrix of the generated global surrogate graph by Eq. 9. We implement the as a simple 3-layer MLP with input dimensions ( is the node feature dimension) and output dimension is 1 (output , represents the probability of existence of a link). Here we compare the other two different implementations of : (1) make a learnable matrix and directly optimize , (2) is determined by the dot product of (normalized) node features. The experimental results are shown in Table 10. The advantages of the other two designs cannot be generalized to most datasets. Our design, in contrast, offers flexibility via the learnable MLP and value, thus showing overall performance advantages. Note that when setting , our design falls back to derive a global surrogate graph where all nodes are isolated.
| Cora | CiteSeer | PubMed | ogbn-arxiv | |||||
| Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | |
D.5 Performance under Model Heterogeneity
We conduct experiments on 10 clients with heterogeneous models. The model configuration is detailed in Section C. Under model heterogeneity, most pFGL methods (except FedTAD and FedSpray) do not work since they rely on the model parameter aggregation. Thus, we compare our method with FedTAD, FedSpray, DENSE, and Co-Boost. The results under Louvain and Metis partitions are presented in Table 11 and Table 12. The results demonstrate the superior performance of our method in both accuracy and F1-macro.
| Method | Cora | CiteSeer | ogbn-arxiv | |||
| Acc. | F1 | Acc. | F1 | Acc. | F1 | |
| O-pFGL | 75.040.50 | 57.261.93 | 70.870.21 | 56.670.20 | 70.980.09 | 36.410.43 |
| Method | Cora | CiteSeer | ogbn-arxiv | |||
| Acc. | F1 | Acc. | F1 | Acc. | F1 | |
| O-pFGL | 79.640.94 | 44.862.04 | 71.820.34 | 47.570.98 | 70.310.06 | 40.310.11 |